Please help!
Solve the following equation for x:
10x^2=9x+9
thanks,
Lackinmath

Answer by stanbon(75887)

10x^2=9x+9
Rewrite as:
10x^2-9x-9=0
Use the AC Method to Rewrite as:
10x^2-15x+6x-9=0
Factor the 1st two then the last two terms to get:
5x(2x-3)+3(2x-3)=0
Factor again to get:
(2x-3)(5x+3)=0
Set each factor equal to zero and solve:
x=3/2 or x=-3/5
Cheers,
Stan H.